Solve the equation 12x^2 - 23x + 8 = -2
Write in set notation.

Answer by richwmiller(17219)

12x^2 - 23x + 8 = -2
12x^2 - 23x + 10 = 0
find factors of 120 which add up to -23
First Number Second Number Sum
1 120 1+120=121
2 60 2+60=62
3 40 3+40=43
4 30 4+30=34
5 24 5+24=29
6 20 6+20=26
8 15 8+15=23
10 12 10+12=22
-1 -120 -1+(-120)=-121
-2 -60 -2+(-60)=-62
-3 -40 -3+(-40)=-43
-4 -30 -4+(-30)=-34
-5 -24 -5+(-24)=-29
-6 -20 -6+(-20)=-26
-8 -15 -8+(-15)=-23
-10 -12 -10+(-12)=-22
-8 and -15
12x^2-8x-15x+10=0
4x*(3x-2)-5*(3x-2)=0
(4x-5)*(3x-2)=0
4x-5=0
4x=5
x=5/4
3x-2=0
3x=2
x = 2/3
(2/3,5/4)